### v1

@@ 5.1 What do Attractive Employers Offer in Job Ads?
Are attractive employers more likely to mention certain job attributes in publicly posted vacancies? We first investigate whether employers estimated as high-pay or low-pay advertise different types of attributes in their job postings. We estimate the regression model:
@@
at the establishment-cluster level $g$, where $\text{Attribute}_{g(j)}^{k}$ is the share of job ads from establishments in cluster $g$ that advertise attribute $k$, $\psi_{g(j)}$ is their estimated pay component, $X_{g(j)}$ is a vector of occupation, industry, and location controls, and $\varepsilon_{g(j)}^{k}$ is an error term. [^41] For each establishment-cluster g, the vector $X_{g(j)}$ controls flexibly for the composition of workers and establishments in our matched employer-employee data, with respect to 2-digit industries, 2-digit occupations, and 10 location groups. The patterns we describe below are robust to dropping controls $X_{g(j)}$. The coefficient of interest is $\alpha_{\psi}^{k}$, which measures to what extent employers with a standard deviation higher pay premium tend to advertise job attribute $k$ conditional on the controls in $X_{g(j)}$. We weight by the number of worker-years in each establishment-cluster.
We proceed similarly to study which attributes employers estimated as having high amenity value are more likely to advertise, substituting $a_{g(j)}$ for $\psi_{g(j)}$ in Equation (7), from which we obtain a second set of coefficients of interest $\alpha_{a}^{k}$ for each advertised job attribute $k$. While the pay and amenity estimates are negatively correlated, as shown in Panel C of Figure 4, there is substantial variation in the employer non-pay values also conditional on the employer pay premiums, and therefore, $\alpha_{a}^{k}$ is not proportional to $-\alpha_{\psi}^{k}$.
@@
Note: This figure shows parameter estimates from separate regressions of the share of each ad attribute on employer pay ($\psi_{j}$) and non-pay ($a_{j}$) values, capturing the change in the fraction of ads with each job attribute associated with a standard deviation increase in employer value. Estimations are done at the employer-cluster level for the five-year period from 2015 to 2019, and are weighted by the number of worker-years. We control for occupation and industry using two-digit occupation and two-digit industry fixed effects and location based on deciles of the number of workers in the municipality of the firm (Oslo as an own group). Appendix Figure A.5 shows corresponding results by overall employer values ($V_{j}$).
@@
Figure 5 reports the estimated $\alpha_{\psi}^{k}$ and $\alpha_{a}^{k}$ for all attributes we derive from the vacancy data. Panel (a) shows that high-pay employers are more likely to mention characteristics associated with a more generous compensation package (“pension/insurance scheme,” “competitive pay,” “collective agreement pay”) and career opportunities (“career paths,” “on-the-job training”). Perhaps more surprisingly, we also find that several attributes related to the workplace, such as “good colleagues” and “social environment,” are associated with high-pay employers. At the other end of the spectrum, we see that most low-pay employers are more likely to specify the type of hours and the type of contract in their job postings. Panel (b) further shows that high non-pay employers are more likely to specify working hours and contract duration in the text of their job vacancies, especially in relation to a predictable schedule (“regular daytime schedule”) or flexible schedule (“possibility to work flexible hours”). Conversely, several career-related (“on-the-job training”) and workplace-related (“social environment,” “good colleagues”) attributes are negatively associated with non-pay values. Non-standard hours (“shift work”) is also more frequently advertised by employer with low non-pay value. Finally, Figure 5 shows that many estimated $\alpha_{\psi}^{k}$ and $\alpha_{a}^{k}$ are not statistically different from zero. We therefore fail to detect any correlation between many advertised attributes and the estimated value of the posting employers.
@@
Overall, the correlations that we document in Figure 5 are consistent with the existence of compensating differentials, in line with the negative correlation between the pay and non-pay model estimates (Panel E in Figure 4). Notably, jobs offered by employers with higher non-pay value appear to more frequently have convenient hours, echoing the findings obtained in experimental settings (Mas and Pallais, 2017). However, the more accurate interpretation of the coefficients $\alpha_{\psi}^{k}$ and $\alpha_{a}^{k}$ is as measures of what high-pay and high non-pay employers are more likely to advertise in their ads. The advertised attributes need not translate directly into attributes of the actual jobs for the following reasons. First, it is unclear to what extent employers can deliver on many of these attributes, especially given that some of them are vague (“social environment,” “good colleagues”). Second, employers can omit specific characteristics from the text of the ad even though they are actual attributes of the position. As an example, they may actually offer a competitive salary without mentioning it explicitly in the corresponding vacancy text. As such, we are cautious in interpreting the pay and non-pay attributes mentioned in job ads as reflecting actual job attributes in general.
@@ 5.2 Predicting the Value of Employers from Job Ads
Note: This figure shows adjusted $R^{2}$s from regressions of estimated model parameters on pay and non-pay job attributes and other ad characteristics retrieved in the text analysis, including the number of words in the ad, indicators for ad posted online, disclosure of company name, and the overall number of attributes in the ad. Estimations are done at the employer-cluster level for the five-year period from 2015 to 2019, and are weighted by the number of worker-years in each cluster. Overall $R^{2}$s are from regressions that include ad attributes and other ad characteristics, while partial $R^{2}$s are from regressions that also control for the composition of industry, occupation and location in each cluster. Industry and occupation controls are defined at the two-digit level. Location indicators split local municipalities into deciles, with municipalities in the same group having similar number of workers (Oslo as an own group). Appendix Figure A.7 provides corresponding results separately for each of the four five-year periods between 2000 and 2019.
Summarizing the overall predictive power job ads, Figure 6 shows the adjusted-$R^{2}$ of the regression model (8) with and without controlling for the variables in $X_{g}.$ [^42] The formula used to calculate predictive power is as follows: Adjusted $\textnormal{Adjusted }R^{2}=1-\frac{SS_{res}/df_{res}}{SS_{total}/df_{total}}$ where $SS_{res}$ is residual sum of squares, $df_{res}$ is the number of establishment clusters (approximately 1,560 in our baseline) minus the number of explanatory variables, and $SS_{total}$ and $df_{total}$ are from regressions on a constant term only. The partial $R^{2}$ is equal to the share of the variation that is unexplained in a regression on baseline controls only but explained when data extracted from job ads is added to the right hand side: Partial $\textnormal{Partial }R^{2}=\frac{R^{2}_{full}-R^{2}_{baseline}}{1-R^{2}_{baseline}}$ where $R^{2}_{full}$ ($R^{2}_{baseline}$) is adjusted-$R^{2}$ from a regression on baseline controls and (without) job attributes. This plot shows that the information contained in job ads meaningfully predicts overall employer values and their components. We find that the non-pay attributes (blue bars) are about as predictive of the overall employer values as the pay-related attributes (red bars). In total, solely based on the job attributes we are able to detect in ads (i.e., without including the controls in $X_{g(j)}$), we are able to explain about 60 percent of the variation in employer pay and their overall values, and about 50 percent of non-pay and 40 percent of the variation in job security. Once we purge the information contained in $X_{g}$, we find that detected job attributes still explain around 10 to 20 percent of the remaining variation in employer values, as shown by the partial-$R^{2}$ in Figure 6.
@@
We further decompose the adjusted-$R^{2}$ measures reported above into contributions from different categories of job attributes in Appendix Figure A.6. Panel (a) documents that financial attributes explain a large share of the variation in employer pay values, while hours of work and convenient hours contain some of the residual variation. Similarly, Panel (b) shows that a large fraction of the variation in non-pay employer values can be explained by detected job attributes related to convenient and inconvenient hours, while pay-related attributes contribute less to explanatory power. In a similar spirit, in Appendix D, we further document that the patterns discussed here are robust to focusing on specific job attributes and the order in which we add various attributes. Specifically, we show evidence from a series of regressions where we sequentially add the individual job attributes that provide the largest increase in predictive power. Based on this approach, we confirm that a substantial share of the predictive power with respect to employer pay values is contained in the information related to financial attributes, while the information on convenient or inconvenient hours remains key in the predictions of employer non-pay values.
@@
Next, we consider how the posted employer values that are predicted using Equation (8) relate to the average duration of vacancies posted by the corresponding employers. Using data on job ads, recent evidence by Mueller et al. (2024) and Bassier et al. (2023) shows that employer pay premiums and posted pay are negatively related to the duration of posted vacancies, consistent with the notion that high-pay employers are more attractive in the labor market. In Appendix Table A.5, Column (1), we show that the average vacancy duration is 15.4 percent shorter for job ads with a standard deviation higher overall posted employer value, while controlling flexibly for industries, occupations and locations listed in the job ads. Decomposing the overall posted employer value in its various components, we find a vacancy duration elasticity of $-0.4$ with respect to the posted pay premium and an elasticity of $-0.5$ with respect to the posted amenity value. Further, we get a negative relationship between posted job security and vacancy duration, finding 8.6 percent shorter average duration for job ads with a standard deviation higher job security value. These results confirm that both pay premiums, amenity values and job security shape the overall attractiveness of employers.
@@
Returning to the regression model in Equation (8), Figure 7 shows the adjusted-$R^{2}$ by gender and by education, with and without controlling for the variables in $X_{g(j)}$. The outcomes in these regressions–the employer values–are estimated separately for each sub-sample, and all regressions are weighted by the corresponding number of worker-years in each sub-sample. These plots show that the information contained in vacancy texts meaningfully predicts the value of employers in all sub-samples. Across sub-samples, the magnitude of the overall $R^{2}$s are broadly similar to those reported for the pooled sample in Figure 6, i.e., around 60 percent for the estimated overall employer values. The partial $R^{2}$s tend to be slightly larger when we split the sample by gender and by education. It is consistently greater than 20 percent for the overall employer values as compared to 15 percent in the pooled sample.
@@
Note: This figure shows adjusted $R^{2}$s from regressions of estimated population group-specific model parameters on pay and non-pay job attributes and other ad characteristics retrieved in the text analysis, including the number of words in the ad, indicators for ad posted online, disclosure of company name, and the overall number of attributes in the ad. Estimations are done at the employer-cluster level for the five-year period from 2015 to 2019, separately by gender (Panels (a)-(b)) and education (Panels (c)-(d)), and are weighted by the number of worker-years in each cluster and population group. Overall $R^{2}$s are from regressions that include ad attributes and other ad characteristics, while partial $R^{2}$s are from regressions that also control for the composition of industry, occupation, and location in each cluster. Industry and occupation controls are defined at the two digit level. Location indicators split local municipalities into deciles, with municipalities in the same group having similar number of workers (Oslo as an own group). Appendix Figure A.8 shows a corresponding plot with weights fixed across this sample split using the overall number of worker-years.
Overall, Figure 7 suggests that the most salient difference across sub-samples in terms of the explanatory power of the content of job ads is found between college and non-college workers. In particular, the explanatory power of ads is at least 10 percentage points lower for the estimated overall and non-pay employer values for non-college workers relative to college workers, both for the overall and partial $R^{2}$. By comparison, the overall $R^{2}$ suggests that the content of job ads has a higher predictive power in terms of employers’ non-pay value for women than for men. However, this result disappears once we control for $X_{g(j)}$. In Appendix Table A.5, Columns (2)-(5), we report the elasticity of vacancy duration with respect to posted pay and amenity values for each sub-sample, finding broadly similar patterns.
@@ 5.3 A Model-Based Measure of Ad Informativeness
Relative Change in Job-to-Job Mobility (in %):
Pay Attributes Pay and Non-Pay Attributes Full Model
@@
Change Relative to Benchmark Scenario:
Pooled sample 0.2 0.7 -0.5 0.6 1.6 -1.0 1.7 6.7 -5.0
@@
In Table 5, we investigate how job-to-job mobility changes under alternative information sets used to predict overall employer values. All results are shown relative to the benchmark linear prediction model in Equation (8) with occupation, industry, and location controls, but no information from vacancy texts. The numbers reported in the top row suggest that having detailed information on job attributes from job ads would increase (reduce) inflows of workers to better (worse) jobs by 1.6 (1.0) percent, as compared to the benchmark scenario where only “one-liner” vacancies listing occupation, industry, and location are posted. This suggests that publicly advertised information on both pay and non-pay attributes can have important implications for the strength and direction of job-to-job flows. Using our full model, we find that upward mobility goes up by 6.7 percent relative to the benchmark. The corresponding figure for downward mobility is 5 percent. This result suggests that publicly advertised information on job attributes explains around 25 percent of the upward mobility in job-to-job flows that we observe in our model. Using this measure of ad informativeness, we also find evidence of heterogeneity by gender and by education. The heterogeneity results in Table 5 imply that the information on job attributes in ads explains more of the upward mobility in job-to-job flows for women than for men (30 percent vs 20 percent) and for college graduates than for non-college workers (35 percent vs 25 percent).
In Table 6, we extend the analysis in Table 5 to upward and downward mobility in employer pay and non-pay values. Using the same linear models to define counterfactual mobility flows, we now quantify to what extent detailed ad information translates into additional mobility toward better-paying employers $\sum_{j,k}\widehat{M}_{jk}\cdot\mathbf{1}\{\psi_{k}\geq\psi_{j}\}$ or employers offering better amenities $\sum_{j,k}\widehat{M}_{jk}\cdot\mathbf{1}\{a_{k}\geq a_{j}\}$. We similarly define downward mobility in pay and non-pay.
@@
Relative Change in Job-to-Job Mobility (in %):
Pay Pay and Non-Pay
Attributes Attributes Full Model
@@
Change Relative to Benchmark Scenario:
Pooled sample
   Pay 0.5 -0.2 0.4 0.3 0.7 0.1
   Non-pay -0.1 0.3 0.5 0.1 2.1 -0.4
@@
   College
    Pay 0.1 0.1 0.2 0.8 1.0 1.1
    Non-pay 0.4 -0.3 1.6 -0.6 2.7 -0.6
   Non-college
    Pay 1.3 -0.6 1.4 -0.4 1.8 0.5
    Non-pay -0.1 0.7 0.3 0.8 1.7 0.6
@@
Note: The table reports relative changes in job-to-job mobility rates (in %) under alternative information sets used to predict the overall employer values, as compared to a benchmark scenario where workers can only use a linear projection based on industry, occupation and location. Each row reports upward and downward mobility in Pay ($\psi_{j}$) or Non-pay ($a_{j}$) for each sub-sample. The columns denoted by $\uparrow$ shows the rate of change in upward mobility in Pay ($\psi_{k}$geq $\psi_{j}$) and Non-pay ($a_{k}$geq $a_{j}$), respectively. The columns denoted by $\downarrow$ shows the rate of change in downward mobility in Pay ($\psi_{k}$ $<$ $\psi_{j}$) and Non-pay ($a_{k}$ $<$ $a_{j}$), respectively.
@@
We draw two main conclusions from this exercise. First, we find that the magnitude of the changes in job-to-job flows to employers offering better pay and better non-pay is smaller across prediction models, as compared to Table 5. For instance, the increase in job-to-job flows to employers with higher value is 1.6 percent using the predictive model with both pay and non-pay attributes in the pooled sample (column (5) top row in Table 5). This number drops to 0.4 percent (0.5 percent) for flows to employers offering better pay (non-pay). This reflects the relatively limited correlation between the overall value of employers $V_{j}$ and its components: the correlation between $\psi_{j}$ and $V_{j}$ is 0.36 in the pooled sample (see Table 4). Second, this exercise suggests that there is heterogeneity in the direction of job-to-job flows predicted by the pay and non-pay attributes contained in vacancy texts. Focusing on gender differences, we find that men are more likely to move to employers offering better pay as their information set is expanded. Relative to the baseline model, their job-to-job mobility to better paying employers increases by 0.5 percent using the full set of attributes advertised in vacancies (Column (3) in Table 6). By contrast, women’s job-to-job mobility to employers offering higher pay decreases with the full set of advertised attributes. The flip side of this decrease in job-to-job flows to higher-paying employers is a one percent increase in the flows to employers offering better amenities. The intuition behind the gender results is straightforward. In the estimated model, pay is a more important component of the value of employers for men than for women, as shown in Table 4. We also find that advertised attributes are slightly more informative of the value of employers for women than for men (Table 5). As a result, with more information, women self-select more into high-amenity jobs. This last finding is consistent with gender differences in preferences over employer attributes carrying over to the job search stage (e.g., Le Barbanchon et al., 2021; Fluchtmann et al., 2024).

### v2

@@ 5.1 What do Attractive Employers Offer in Job Ads?
Are attractive employers more likely to mention certain job attributes in publicly posted vacancies? We first investigate whether employers estimated as high-pay or low-pay advertise different types of attributes in their job postings. We estimate the regression model
@@
at the establishment-cluster level $g$, where $\text{Attribute}_{g(j)}^{k}$ is the share of job ads from establishments in cluster $g$ that advertise attribute $k$, $\psi_{g(j)}$ is their estimated pay component, $X_{g(j)}$ is a vector of occupation, industry, and location controls, and $\varepsilon_{g(j)}^{k}$ is an error term. The coefficient of interest is $\alpha_{\psi}^{k}$, which measures to what extent employers with one standard deviation higher pay premium tend to advertise job attribute $k$ conditional on the controls in $X_{g(j)}$. [^38] For each establishment-cluster g, the vector $X_{g(j)}$ controls flexibly for the composition of workers and establishments in our matched employer-employee data, with respect to 2-digit industries, 2-digit occupations, and 10 location groups. The patterns we describe below are robust to dropping controls $X_{g(j)}$. We weight by the number of worker-years in each establishment-cluster.
We proceed similarly to study which attributes employers estimated as having high amenity value are more likely to advertise, substituting $a_{g(j)}$ for $\psi_{g(j)}$ in Equation (7), from which we obtain a second set of coefficients of interest $\alpha_{a}^{k}$ for each advertised job attribute $k$. While the pay and amenity estimates are negatively correlated, as shown in Panel D of Figure 4, there is substantial variation in the employer non-pay values also conditional on the employer pay premiums, and therefore, $\alpha_{a}^{k}$ is not mechanically equal to $-\alpha_{\psi}^{k}$.
@@
Note: This figure shows parameter estimates and $90\%$ confidence intervals from separate regressions of the share of each ad attribute on employer pay ($\psi_{j}$) and non-pay ($a_{j}$) values, capturing the change in the fraction of ads with each job attribute associated with a standard deviation increase in employer value. Estimations are done at the employer-cluster level for the five-year period from 2015 to 2019, and are weighted by the number of worker-years. We control for occupation and industry using two-digit occupation and two-digit industry fixed effects and location based on deciles of the number of workers in the municipality of the firm (Oslo as an own group). Appendix Figure A.5 shows corresponding results by overall employer values ($V_{j}$) for the baseline model specification, while Appendix Figures A.6-A.7 shows results for model estimates using smaller and larger employer clusters, respectively.
@@
Figure 5 reports the estimated $\alpha_{\psi}^{k}$ and $\alpha_{a}^{k}$ for all attributes we derive from the vacancy data. Panel (a) shows that high-pay employers are more likely to mention characteristics associated with a more generous compensation package (“pension/insurance scheme,” “competitive pay,” “collective agreement pay”) and career opportunities (“career paths,” “on-the-job training”). Perhaps more surprisingly, we also find that several attributes related to the workplace, such as “good colleagues” and “social environment,” are associated with high-pay employers. At the other end of the spectrum, we see that most low-pay employers are more likely to specify the type of hours and the type of contract in their job postings. Panel (b) further shows that high amenity employers are more likely to specify contract duration and hours-related attributes in the text of their job vacancies, especially in relation to a predictable schedule (“regular daytime schedule”) or flexible schedule (“possibility to work flexible hours”). Conversely, several career-related (“on-the-job training”) and workplace-related (“social environment,” “good colleagues”) attributes are negatively associated with amenity values. Non-standard hours (“shift work”) is also more frequently advertised by employer with low amenity value. Finally, Figure 5 shows that many estimated $\alpha_{\psi}^{k}$ and $\alpha_{a}^{k}$ are not statistically different from zero. We therefore fail to detect any correlation between many advertised attributes and the estimated value of the posting employers.393939 Appendix Figure A.5 shows results corresponding to Figure 5 for the overall employer value $V_{j}$, while Appendix Figures A.6A.7 show results, respectively, using smaller and larger employer-clusters.
@@
Overall, the correlations that we document in Figure 5 are consistent with the existence of compensating differentials, in line with the negative correlation between the pay and amenity model estimates (Panel D in Figure 4). Notably, jobs offered by employers with higher non-pay value appear to more frequently have convenient hours, echoing the findings obtained in experimental settings (Mas and Pallais, 2017). However, a more accurate interpretation of the coefficients $\alpha_{\psi}^{k}$ and $\alpha_{a}^{k}$ is as measures of what high-pay and high non-pay employers are more likely to advertise in their ads. The advertised attributes need not translate directly into attributes of the actual jobs for the following reasons. First, it is unclear to what extent employers can deliver on many of these attributes, especially given that some of them are vague (“social environment,” “good colleagues”). Second, employers can omit specific characteristics from the text of the ad even though they are actual attributes of the position. As an example, they may actually offer a competitive salary without mentioning it explicitly in the corresponding vacancy text. As such, we are cautious in interpreting the pay and non-pay attributes mentioned in job ads as reflecting actual job attributes in general.
@@ 5.2 Predicting the Value of Employers from Job Ads
Summarizing the overall predictive power of job ads, Figure 6 shows the adjusted-$R^{2}$ of the regression model (8) with and without controlling for the variables in $X_{g}.$ [^40] The formula used to calculate predictive power is as follows: Adjusted $\textnormal{Adjusted }R^{2}=1-\frac{SS_{res}/df_{res}}{SS_{total}/df_{total}}$ where $SS_{res}$ is residual sum of squares, $df_{res}$ is the number of establishment clusters (approximately 1,560 in our baseline) minus the number of explanatory variables, and $SS_{total}$ and $df_{total}$ are from regressions on a constant term only. The partial $R^{2}$ is equal to the share of the variation that is unexplained in a regression on baseline controls only but explained when data extracted from job ads is added to the right hand side: Partial $\textnormal{Partial }R^{2}=\frac{R^{2}_{full}-R^{2}_{baseline}}{1-R^{2}_{baseline}}$ where $R^{2}_{full}$ ($R^{2}_{baseline}$) is adjusted-$R^{2}$ from a regression on baseline controls and (without) job attributes. This plot shows that the information contained in job ads meaningfully predicts overall employer values and their components. We find that the non-pay attributes (blue bars) are about as predictive of the overall employer values as the pay-related attributes (red bars). In total, solely based on the job attributes we are able to detect in ads (i.e., without including the controls in $X_{g(j)}$), we are able to explain about 60 percent of the variation in employer pay and their overall values, and about 50 percent of non-pay and 40 percent of the variation in job security. Once we purge the information contained in $X_{g}$, we find that detected job attributes still explain around 10 to 20 percent of the remaining variation in employer values, as shown by the partial-$R^{2}$ in Figure 6.414141 Appendix Figure A.10 shows that alternatively using smaller or larger employer-clusters yields the same overall pattern in terms of the predictive power of ad attributes for each component of the value of employers.
@@
Note: This figure shows adjusted $R^{2}$s from regressions of estimated model parameters on pay and non-pay job attributes and other ad characteristics retrieved in the text analysis, including the number of words in the ad, indicators for ad posted online, disclosure of company name, and the overall number of attributes in the ad. Estimations are done at the employer-cluster level for the five-year period from 2015 to 2019, and are weighted by the number of worker-years in each cluster. Overall $R^{2}$s are from regressions that include ad attributes and other ad characteristics, while partial $R^{2}$s are from regressions that also control for the composition of industry, occupation and location in each cluster. Industry and occupation controls are defined at the two-digit level. Location indicators split local municipalities into deciles, with municipalities in the same group having similar number of workers (Oslo as an own group). The sample corresponds to the last row in Table 1. Appendix Figure A.10 reports the corresponding results using employer values estimated on larger and smaller clusters, while Appendix Figure A.11 provides results separately for each of the four five-year periods between 2000 and 2019.
@@
We further delve into the adjusted-$R^{2}$ measures reported in Figure 6 with the following two exercises. First, we decompose the adjusted-$R^{2}$ measures reported above into contributions from different categories of job attributes in Appendix Figure A.8. Panel (a) documents that financial attributes explain a large share of the variation in employer pay values, while hours of work and convenient hours contain some of the residual variation. Similarly, Panel (b) shows that a large fraction of the variation in non-pay employer values can be explained by detected job attributes related to convenient and inconvenient hours, while pay-related attributes contribute less to explanatory power. In a similar spirit, in Appendix D, we further document that the patterns discussed here are robust to focusing on specific job attributes and the order in which we add various attributes. Specifically, we show evidence from a series of regressions where we sequentially add the individual job attributes that provide the largest increase in predictive power. Based on this approach, we confirm that a substantial share of the predictive power with respect to employer pay values is contained in the information related to financial attributes, while the information on convenient or inconvenient hours remains key in the predictions of employer non-pay values.
Second, we consider the predictive power of job ads across different industries in Appendix Figure A.9. We estimate the regression models in Equation (8) separately by industry. [^42] We assign an industry to each cluster $g$ based on modal industry among the establishments in the cluster. While job ads hold strong predictive power for the overall values of employers across all industry groups, we do find sizeable differences across industries. Notably, the adjusted-$R^{2}$ ranges from 0.5 in the Construction industry to 0.8 in the Retail and Services industry.
Posted Employer Value and Vacancy Duration
@@
Next, we consider how the employers’ values posted in job ads–as can be predicted using the regression model in Equation (8)–relate to the average duration of vacancies that these employers open. Recent evidence provided by Mueller et al. (2024) and Bassier et al. (2023), respectively, shows that the employer pay premiums and the posted pay are negatively related to the duration of posted vacancies, consistent with the notion that high-pay employers are more attractive in the labor market. In Appendix Table A.4, we show that the average vacancy duration is 13.9 percent shorter for job ads with one standard deviation higher overall posted employer value, while controlling flexibly for industries, occupations and locations listed in the job ads. Decomposing the overall posted employer value in its various components, we find a vacancy duration elasticity of $-0.7$ with respect to the posted pay premium and an elasticity of about $-1$ with respect to the posted amenity value. Further, we get a negative relationship between posted job security and vacancy duration, finding 5.9 percent shorter average duration for job ads with a standard deviation higher job security value. These results show that both pay premiums, amenity values and job security shape the overall attractiveness of employers.
@@
Returning to the regression model in Equation (8), Figure 7 shows the adjusted-$R^{2}$ by gender and by education, with and without controlling for $X_{g(j)}$. The outcome variables in these regressions–the employer values–are estimated separately for each sub-sample, and all regressions are weighted by the corresponding number of worker-years in each sub-sample. These plots show that the information contained in vacancy texts meaningfully predicts the value of employers in all sub-samples. Across sub-samples, the magnitude of the overall $R^{2}$s is around 60 percent for the estimated overall employer value, broadly similar to those reported for the pooled sample in Figure 6.
@@
Note: This figure shows adjusted $R^{2}$s from regressions of estimated population group-specific model parameters on pay and non-pay job attributes and other ad characteristics retrieved in the text analysis, including the number of words in the ad, indicators for ad posted online, disclosure of company name, and the overall number of attributes in the ad. Estimations are done at the employer-cluster level for the five-year period from 2015 to 2019, separately by gender (Panels (a)-(b)) and education (Panels (c)-(d)), and are weighted by the number of worker-years in each cluster and population group. Overall $R^{2}$s are from regressions that include ad attributes and other ad characteristics, while partial $R^{2}$s are from regressions that also control for the composition of industry, occupation, and location in each cluster. Industry and occupation controls are defined at the two digit level. Location indicators split local municipalities into deciles, with municipalities in the same group having similar number of workers (Oslo as an own group). Appendix Figure A.12 shows a corresponding plot with weights fixed across sample splits using the total number of worker-years.
Overall, Figure 7 suggests that the most salient difference across sub-samples in terms of the explanatory power of the content of job ads is found between college and non-college workers. In particular, the explanatory power of ads is at least 10 percentage points lower for the estimated overall and non-pay employer values for non-college workers relative to college workers, both for the overall and partial $R^{2}$. By comparison, the overall $R^{2}$ suggests that the content of job ads has a higher predictive power in terms of employers’ non-pay value for women than for men. However, this result disappears once we control for $X_{g(j)}$. [^43] Appendix Figure A.11 provides similar heterogeneity results over time, separately for each of the four five-year periods between 2000 and 2019. These results do not suggest a clear time profile.
@@ 5.3 A Model-Based Measure of Ad Informativeness
Relative Change in Job-to-Job Mobility (%)
Pay Pay & Non-Pay Full Model
@@
% change relative to baseline scenario:
Pooled sample 0.1 0.5 -0.3 0.3 1.1 -0.8 1.1 4.3 -3.2
@@
In Table 5, we investigate how job-to-job mobility changes under alternative information sets used to predict overall employer values. All results are shown relative to the benchmark linear prediction model in Equation (8) with occupation, industry, and location controls, but no information from vacancy texts. The numbers reported in the top row suggest that having detailed information on job attributes from job ads would increase (reduce) inflows of workers to better (worse) jobs by 1.1 (0.8) percent, as compared to the benchmark scenario where only “one-liner” vacancies listing occupation, industry, and location are posted. This suggests that publicly advertised information on both pay and non-pay attributes can have important implications for the strength and direction of job-to-job flows. Using our full model, we find that upward mobility goes up by 4.3 percent relative to the benchmark. The corresponding figure for downward mobility is 3.2 percent. This result suggests that publicly advertised information on job attributes explains around 25 percent of the upward mobility in job-to-job flows that we observe in our model. Using this measure of ad informativeness, we also find evidence of heterogeneity by gender and by education. The heterogeneity results in Table 5 imply that the information on job attributes in ads explains slightly more of the upward mobility in job-to-job flows for women than for men (30 percent vs 20 percent) and for college graduates than for non-college workers (35 percent vs 25 percent).
In Table 6, we extend the analysis to upward and downward mobility in employer pay and amenity values. Using the same linear models to define counterfactual mobility flows, we now quantify to what extent detailed ad information translates into additional mobility toward better-paying employers $\sum_{j,k}\widehat{M}_{jk}\cdot\mathbf{1}\{\psi_{k}\geq\psi_{j}\}$ or employers offering better amenities $\sum_{j,k}\widehat{M}_{jk}\cdot\mathbf{1}\{a_{k}\geq a_{j}\}$. We similarly define downward mobility in pay and non-pay.
@@
Relative Change in Job-to-Job Mobility (%)
Pay Pay & Non-Pay Full Model
@@
Panel A: Pay mobility
% change relative to baseline scenario:
Pooled sample 0.3 -0.1 0.2 0.1 0.5 0.7
@@
   College -0.0 0.1 -0.1 0.5 0.3 0.6
   Non-college 0.8 -0.5 0.9 -0.3 1.4 0.2
Panel B: Non-pay mobility
% change relative to baseline scenario:
Pooled sample 0.1 0.0 0.6 -0.3 2.6 -1.5
Gender:
   Women 0.3 -0.3 1.1 -0.8 2.2 -1.5
   Men 0.0 0.1 0.4 -0.1 2.3 -1.0
Education:
   College 0.5 -0.4 1.3 -0.9 2.7 -1.8
   Non-college -0.1 0.4 0.3 0.3 2.0 -0.4
@@
Note: The table reports relative changes in job-to-job mobility rates (in %) under alternative information sets used to predict the overall employer values, as compared to a benchmark scenario where workers can only use a linear projection based on industry, occupation, and location. Panel A reports upward and downward mobility in Pay ($\psi_{j}$) for each sub-sample. Panel B reports upward and downward mobility in Non-pay ($a_{j}$) for each sub-sample. The columns denoted by $\uparrow$ shows the rate of change in upward mobility in Pay ($\psi_{k}$geq $\psi_{j}$) and Non-pay ($a_{k}$geq $a_{j}$), respectively. The columns denoted by $\downarrow$ shows the rate of change in downward mobility in Pay ($\psi_{k}$ $<$ $\psi_{j}$) and Non-pay ($a_{k}$ $<$ $a_{j}$), respectively.
@@
We draw two main conclusions from this exercise. First, we find that the magnitude of the changes in job-to-job flows to employers offering better pay and better non-pay is smaller across prediction models, as compared to Table 5. For instance, in the top row of Table 5, Column (5), the increase in job-to-job flows to employers with higher value is 1.1 percent using the predictive model with both pay and non-pay attributes in the pooled sample. As shown in Table 6, Column (3), job-to-job flows to employers offering better pay (non-pay) increase by only 0.2 percent (0.6 percent). These results reflect the relatively limited correlation between the overall value of employers $V_{j}$ and its components: the correlation between $\psi_{j}$ and $V_{j}$ is 0.36 in the pooled sample; see Table 4. Second, this exercise suggests that there is heterogeneity in the direction of job-to-job flows predicted by the pay and non-pay attributes contained in vacancy texts. Focusing on gender differences, we find that men are more likely to move to employers offering better pay as their information set is expanded. Relative to the baseline model, their job-to-job mobility to better paying employers increases by 0.3 percent using the full set of attributes advertised in vacancies (Panel A, Column (3) in Table 6). By contrast, women’s job-to-job mobility to employers offering higher pay decreases with the full set of advertised attributes. The flip side of this decrease in job-to-job flows to higher-paying employers is a one percent increase in the flows to employers offering better amenities. The intuition behind the gender results is straightforward. In the estimated model, pay is a more important component of the value of employers for men than for women, as shown in Table 4. We also find that advertised attributes are slightly more informative of the value of employers for women than for men (Table 5). As a result, with more information, women self-select more into high-amenity jobs. This last finding is consistent with gender differences in preferences over employer attributes carrying over to the job search stage (e.g., Le Barbanchon et al., 2021; Fluchtmann et al., 2024).